log2 8=x
x is equivalent to 3 because 2*2*2=8 Is this how you do this?

Found 2 solutions by sarah_adam, Nate:
Answer by sarah_adam(201)

log2 8 = x
log2 (2^3) = x
3 log2 2 = x
log2 to the base 2 = 1
3(1) = x
x = 3

Answer by Nate(3500)

Here is another way that is similiar:
log(base 2) 8 = x
2^x = 8 exponential form
log (2^x) = log (8)
x log (2) = log (8) use product property
x = (log (8))/(log (2))
x = 3 use a calculator